Shift report 20011114 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Nov 14 2001 - 16:20:28 EST

  • Next message: Flemming Videbaek: "Fw: Decision on Running at 22 GeV This Year"

    ------------------------------------------------------------
    Date: Wed Nov 14 16:20:27 2001, Shift: 08:00-16:00
    Supervisor: ykl, with: 
    Log book#10 pages 139 - 140
    
    Shift Summary:
    ------------------------------------------------------------
    1. Machine had cryogenic problems and ramping problems until 14:10.
    Then they had autoprogram program so that steering was delayed until
    15:30. 
    
    2
    Run 5793 with trigger 6 of 51k. MRS at 40 degrees and 4B; 
    FFS at 12 degrees and BFS at 6 degrees with 1/2A.This run
    is short , taken while waiting for a dump. Run 5793 had T3 off.
    
    Run 5797 with trigger 6 of 54k. MRS at 40 degrees and 4B;
    FFS and BFS at 12 degrees with 1/2A. This run is short while
    the beam was being steered.
    
    Fun 5798 is in progress.
    
    Pedestal runs 5794, 5795, 5796 for TPC, Detectors, DVM, respectively.
    
    3. No additional detector problems.
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Nov 14 2001 - 16:20:57 EST